Following Moroi in [?] I see that for large tan β g-2 goes like

δa_{μ} ∝ \frac{m_{μ}^{2}μ tan β}{M_{1/2}} J  (1)

where J is a loop factor depending on μ and M_{1/2} slightly on M_{0}. For large M_{1/2}, J ≃ 1/M_{1/2}

Following Chankowski et.al and Urban et.al [my program] the Wilson coefficients c_{S,P} in B(B̄_{s} → μ^{+}μ^{−}) and for go like

|c_{S,P} | ∝ \frac{m_{μ}m_{t}^{2}μ tan β^{3}}{M_{1/2}}  (2)